a charge is uniformly distributed on a circular ring of radius r now the ring is rotated with angular velocity w about the diameter of the ring. calculate the magnetic moment of the circular ring

Grade:12

2 Answers

Vikas TU
4 years ago
Hiii 
I am calculating the magnetic moment of the circular ring at the centre of the ring.
The total magnetic induction at the centre of the ring will be  mew* q*w/4*pie*r
I = q/t
I = q*w/2*pie
Magnetic field at the centre of current carrying conductor = mew*I / 2*r
Putting the value of I we ger B = mew* q*w/4*pie*r
Khimraj
4 years ago
The total magnetic induction at the center of the ring will be μ₀qω/4πr
Explanation:
since the total charge is q and we know that current = charge/time
=> I = q/t
The angular velocity is equal to ω = 2π/t
Time period, t = 2π/ω
=> I = qω/2π
we know that magnetic field at the center of a current carrying conductor,
B = μ₀I/2r
putting the value of I from the above equation, we get
B = (μ₀/2r) x qω/2π
=> B = μ₀qω/4πr
Hence the total magnetic induction at the center of the ring will be μ₀qω/4πr